OpenStudy (anonymous):

determine whether each infinite geometric series converges or diverges. if the series converges, state the sum. i have to problems 1) 150+30+6+...... 2) -10-20-40-.....

OpenStudy (anonymous):

help @whpalmer4

OpenStudy (anonymous):

second one doesn't converge for sure, the terms do not even go to zero

OpenStudy (anonymous):

ok so there is one how do we find the first one????

OpenStudy (anonymous):

first one has a common ratio of \(r=\frac{1}{5}\) compute \[\frac{150}{1-\frac{1}{5}}\] to add it up

OpenStudy (anonymous):

you can almost do it in your head what is \(1-\frac{1}{5}\) ?

OpenStudy (anonymous):

um .8???

OpenStudy (anonymous):

or 4/5

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

will the sum be 7.5????

OpenStudy (whpalmer4):

\[\frac{150}{1-\frac{1}{5}} = \frac{150}{\frac{4}{5}} = 150 * \frac{5}{4} = \]

OpenStudy (anonymous):

dividing by \(\frac{4}{5}\) is the same as multiplying by \(\frac{5}{4}\) compute \[150\times \frac{5}{4}\]

OpenStudy (anonymous):

ooh, what @whpalmer4 said